OpenStudy (anonymous):

(sqrt5 - sqrt-4)(sqrt35 - sqrt-28)

OpenStudy (john_es):

\[\left(\sqrt{5}-2 i\right) \left(\sqrt{35}-2 i \sqrt{7}\right)=\sqrt{7}-4 i \sqrt{35}\]